OpenStudy (anonymous):

simplify by factoring square root of54

OpenStudy (zpupster):

factors of 54 which is a perfect square?? 1, 2, 3, 6, 9, 18, 27, 54

OpenStudy (zpupster):

\[\sqrt{x}\times \sqrt{y}=\sqrt{xy}\]

jhonyy9 (jhonyy9):

so you know that 54 = 2*27 but you know too that 27=3*9 so now you see that here 9 is a perfect square so than 54= 2*3*3^2 so what mean that sqrt 54 = 3sqrt 6 hope that is understandably and that this will help you

OpenStudy (nirmalnema):



OpenStudy (nirmalnema):